OpenStudy (anonymous):

Simplify. Remember to use absolute value notation when necessary. If root cannot be simplified state that.

OpenStudy (anonymous):

Square root of 8/y^8

OpenStudy (anonymous):

sqrt(8/y^8) \[\sqrt{\frac{ 8 }{ y ^{8} }}\] \[\sqrt{8}=2\sqrt{2}\] and \[\sqrt{y ^{8}}=y^{4}\] So, \[\frac{ 2\sqrt{2} }{ y^{4} }\]